(a+b)=a3+3a^b+3ab^+b3 & (a+b)^=a^+b^+2ab
Answer by ikleyn(53763)

.
(a+b)^3=a3+3a^2b+3ab^2+b^3,
(a+b)^2=a^2+b^2+2ab
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Your post doesn't contain a question, so I don't know which help do you want to get from us.

Also, there are a lot of errors in your writing of formulas.
So, if your goal was to make bad impression about yourself, you almost reached it.

In any case, you can look in the lessons
    - The square of the sum formula
    - The cube of a sum formula
in this site.